# UAV-borne Mapping Algorithms for Canopy-Level and High-Speed Drone Applications

### v1

@@ Abstract
This article presents a comprehensive review of and analysis of state-of-the-art mapping algorithms for UAV (Unmanned Aerial Vehicle) applications, focusing on canopy-level and high-speed scenarios. This article presents a comprehensive exploration of sensor technologies suitable for UAV mapping, assessing their capabilities to provide measurements that meet the requirements of fast UAV mapping. Furthermore, the study conducts extensive experiments in a simulated environment to evaluate the performance of three distinct mapping algorithms: Direct Sparse Odometry (DSO), Stereo DSO (SDSO), and DSO Lite (DSOL). The experiments delve into mapping accuracy and mapping speed, providing valuable insights into the strengths and limitations of each algorithm. The results highlight the versatility and shortcomings of these algorithms in meeting the demands of modern UAV applications. The findings contribute to a nuanced understanding of UAV mapping dynamics, emphasizing their applicability in complex environments and high-speed scenarios. This research not only serves as a benchmark for mapping algorithm comparisons but also offers practical guidance for selecting sensors tailored to specific UAV mapping applications.
@@ 1 Introduction
UAVs, commonly known as drones, have transcended conventional applications to become indispensable tools across an array of disciplines, from environmental monitoring and precision agriculture to disaster response and infrastructure inspection. At the heart of their efficacy lies the sophisticated interplay between UAVs and mapping algorithms, which serve as the backbone for converting raw sensor data into coherent, high-fidelity maps. These algorithms play a pivotal role in navigating complex terrains, extracting meaningful information, and ensuring precise localization of the UAV in real-time. From traditional photogrammetry to advanced techniques like Simultaneous Localization and Mapping (SLAM), mapping algorithms continuously evolve to meet the diverse demands of applications ranging from agriculture and forestry to disaster response and urban planning.
@@
Mapping algorithms for UAVs are significantly influenced by flight altitude, dictating the scale of environmental perception and mapping capabilities. Low-altitude flights excel in detailed and precise mapping, capturing intricate features with high precision. However, as altitude increases, challenges such as reduced sensor performance, diminished feature visibility, and heightened geometric distortions emerge. While 3D or 2D laser scanners generate effective terrain models, their weight and sensitivity to ground proximity pose challenges. Compact depth-sensing devices, though commercially available, often fall short in operational range. Camera-based mapping systems, while lightweight and scalable, face accuracy challenges at high altitudes due to reduced texture and discernible features. This limitation hampers feature tracking and matching, impacting overall mapping algorithm performance. High-altitude flights also amplify drift and uncertainty in UAV trajectory estimation, particularly affecting SLAM algorithms relying on sensor fusion. The accumulation of errors over time compromises pose estimations, emphasizing the critical consideration of flight altitude in optimizing mapping algorithm outcomes.
@@
Mapping algorithms tailored for high-speed UAVs address the specific demands of dynamic and rapid flight scenarios. Real-time operation is imperative, necessitating the synchronization and integration of data from diverse sensors such as LiDAR, cameras, and inertial measurement units (IMUs). Adaptive navigation is equally crucial to accommodate the UAV’s swift maneuvers and maintain mapping precision. Overcoming challenges related to large distances covered between sensor readings during high-speed flights is essential for achieving precise mapping results. Additionally, robustness in the face of environmental variability, including changes in lighting, weather conditions, and terrains, is vital. High-speed UAVs, integral in applications like surveillance and emergency response, benefit from ongoing advancements in mapping algorithms. These improvements enhance effectiveness, allowing UAVs to navigate rapidly changing environments and deliver precise and timely mapping outcomes.
@@
In recent decades, research on mapping algorithms processing image sensor data for 3D geometric reconstructions has thrived, primarily driven by advancements in Structure-from-Motion (SfM) [9, 27, 29, 24, 28, 30, 14, 4, 15] and stereo reconstruction algorithms [26, 18, 8, 2, 7, 6, 12]. These approaches differ in both computation methods and output formats. Structure from Motion analyzes the relative motion of one or multiple cameras through a sequence of 2D images to reconstruct the 3D scene structure. It concurrently tracks camera positions, orientations, and the scene’s 3D structure. In contrast, stereo reconstruction estimates the 3D scene structure using a pair of 2D images captured simultaneously by two cameras with known relative positions. Depth information is derived from disparities in observed scene points between the two images. While both SfM and stereo reconstruction target 3D scene reconstruction, they excel in different applications and scenarios.
This paper aims to contribute more insights into UAV-borne mapping in canopy-level and high-speed drone applications. The contributions of this article include:
@@
  • Conducted a comprehensive analysis, outlining the strengths and limitations of state-of-the-art SfM and stereo reconstruction algorithms.
@@
  • Facilitated benchmarking by comparing the geometry accuracy and computation speed of various mapping algorithms.
@@
  • Presented a theoretical foundation for sensor selection tailored to canopy-level and high-speed UAV mapping applications.
  • Explored the practical implications of mapping algorithms in real-world scenarios, specifically focusing on canopy-level and high-speed drone applications.
@@
  • Introduced an innovative approach to simulate realistic flights, utilizing Unreal Engine for high-realism environment synthesis, Cesium plug-in for geographical context, AirSim for vehicle dynamics, and PX4 Autopilot for precise vehicle control.
This article focuses on surveying existing sensors and advanced reconstruction algorithms for canopy-level and high-speed UAV mapping applications, specifically evaluating Direct Sparse Odometry (DSO) [5], Direct Sparse Odometry with Stereo Cameras (SDSO) [32], and Direct Sparse Odometry Lite (DSOL) [19]. The experimental analysis assesses the mapping performance of these algorithms in terms of speed and accuracy within a simulated environment. By shedding light on the current landscape of these mapping algorithms, this research seeks to propel the capabilities of drones in practical applications where rapid, precise canopy-level mapping is essential.
@@ 2 Related Work
In the realm of 3D reconstruction, this related work section delves into the comparative analysis between Structure-from-Motion and Stereo Reconstruction methodologies. The first subsection elucidates the foundational theories underlying SfM and stereo reconstruction, exploring their distinct characteristics and applications. The second subsection introduces three prominent algorithms central to 3D reconstruction: Direct Sparse Odometry (DSO), Stereo DSO (SDSO), and DSO Lite (DSOL). Notably, DSO embodies the principles of SfM, while SDSO and DSOL leverage stereo reconstruction techniques. This comprehensive overview sets the stage for an in-depth exploration of their strengths, limitations, and practical implications in the ensuing discussions.
2.1 Structure-from-Motion v.s. Stereo Reconstruction
Structure from Motion (SfM) and stereo reconstruction are two leading techniques employed in 3D reconstruction. This subsection meticulously dissects the underlying principles of both methodologies, shedding light on their distinctive attributes and applications.
@@ 2.1.1 Structure-from-Motion
SfM [9] is the process of reconstructing a 3D structure from its projections into a series of images taken from different viewpoints. It leverages the relative movement between a camera and objects in a scene to reconstruct the 3D structure. SfM estimates the camera poses and the spatial arrangement of points in the scene by analyzing the changes in perspective across multiple images. SfM has been extensively studied and applied in diverse fields, including 3D modeling [21, 23], augmented reality [3, 16], autonomous navigation [20, 11], and remote sensing [31, 17]. Researchers have explored various algorithms and optimization methods to enhance the accuracy [27, 29, 24, 28] and efficiency [30, 14, 4, 15] of SfM, making it a robust solution for scenarios where camera poses change dynamically.
@@ 2.1.2 Stereo Reconstruction
Stereo reconstruction involves the process of estimating the 3D structure of a scene from a pair of 2D images captured by two cameras with known relative positions. By analyzing the disparities between the two images, stereo reconstruction algorithms can calculate the depth information of the scene points. This depth information allows for the creation of a 3D representation of the scene. Stereo reconstruction is critical to enabling autonomous capabilities in a wide range of fields including robotics [26, 18], autonomous vehicles [8, 2], and 3D modeling [7, 6, 12].
@@
Fig. 0(a) illustrates the epipolar geometry of two pinhole cameras observing a 3D point $\mathbf{M}$. Stereo reconstruction estimates $\mathbf{M}$’s distance by analyzing its projections $\mathbf{m}$ and $\mathbf{m}^{\prime}$. The baseline $B$ connects camera origins $\mathbf{CL}$ and $\mathbf{CR}$, defining epipolar geometry with epipoles $\mathbf{e}$ and $\mathbf{e}^{\prime}$. The epipolar plane intersects with image planes $\boldsymbol{\pi}$ and $\boldsymbol{\pi}^{\prime}$, forming epipolar lines. According to epipolar geometry, $\mathbf{m}$ in $\boldsymbol{\pi}^{\prime}$ lies on epipolar line $l^{\prime}$. Depth estimation involves finding corresponding points, simplified by image rectification in Fig. 0(b), ensuring $\mathbf{m}$ and $\mathbf{m}^{\prime}$ align. Depth $d$ is estimated through triangulation represented by Eq. 1, considering column differences from $\mathbf{m}$ and $\mathbf{m}^{\prime}$ to the center of the left and right images, baseline $B$, focal length $f$, and pixel width $\delta$ in the rectified image sensor.
$d=\frac{B*f}{\delta*\left(x-x^{\prime}\right)}$ (1)
Figure 2: The dependency between depth estimation accuracy and the baseline of the stereo camera design for a baseline, $\mathbf{B}$ of 34cm, based on [10].
@@ 2.2 Mapping Algorithms
This article focuses on the following three representative state-of-the-art algorithms to investigate their applications to UAV-borne mapping:
  • Structure-from-Motion: Direct Sparse Odometry (DSO) [5]
  • Stereo reconstruction: Direct Sparse Odometry with Stereo Cameras (SDSO) [32] and Direct Sparse Odometry Lite (DSOL) [19]
@@ 2.2.1 DSO: Direct Sparse Odometry
DSO is a visual odometry technique that adapts SfM methods for 3D reconstruction. It directly estimates the camera motion and the sparse 3D structure of the environment from a sequence of 2D images by minimizing photometric errors. DSO differs significantly from traditional techniques by directly optimizing photometric errors in images, without relying on keypoint detectors or geometric priors. For a point, $\mathbf{p}$ in reference frame $I_{i}$, observed as $\mathbf{p^{\prime}}$ in target frame $I_{j}$, the photometric error, given by Eq. 2, is formulated as the weighted Sum of Squared Differences (SSD) over a small neighborhood of pixels.
@@
where $\mathcal{N}_{\mathbf{p}}$ is the set of pixels in the SSD; $(t_{i},t_{j})$ the exposure times of the frame $I_{i}$ and $I_{j}$; $(a_{i},b_{i},a_{j},b_{j})$ the brightness transfer variables defined in DSO for frame $I_{i}$ and $I_{j}$ respectively; And $\|\cdot\|_{\gamma}$ is the Huber norm. In addition to using robust Huber penalties, a gradient-dependent weighting $w_{\mathbf{p}}$ is applied. Further, $\mathbf{p}^{\prime}$ stands for the projected point position of $\mathbf{p}$ with inverse depth $d_{\mathbf{p}}$ ,given by
@@
where $\mathbf{T}_{i},\mathbf{T}_{j}\in\textbf{SE}(3)$ are the camera poses represented by transformation matrices for frame $I_{i}$ and $I_{j}$.
To minimize the photometric error between the corresponding points in two frames, DSO incorporates a fully direct probabilistic model that jointly optimizes all model parameters, including camera motion and geometry, represented as inverse depth in a reference frame. The optimization is accomplished using the Gauss-Newton algorithm in a sliding window [13].
2.2.2 SDSO: Direct Sparse Odometry with Stereo Cameras
SDSO is a stereo version of DSO. In a monocular mapping system like DSO, to initialize the whole system, i.e., to track the second frame with respect to the initial one using Eq. 2, the inverse depth values $d_{p}$ of the points in the first frame are required. In DSO, the points are initialized to have random depth values ranging from 0 to infinity, corresponding to a large depth variance. Unlike that, SDSO uses stereo matching to estimate a semi-dense depth map for the first frame, which significantly increases the tracking accuracy. The constraints from static stereo introduce scale information into the system. They also provide good geometric priors to temporal multi-view stereo.

### v2

# UAV-Borne Mapping Algorithms for Low-Altitude and High-Speed Drone Applications

@@ Abstract
This article presents an analysis of current state-of-the-art sensors and how these sensors work with several mapping algorithms for UAV (Unmanned Aerial Vehicle) applications, focusing on low-altitude and high-speed scenarios. A new experimental construct is created using highly realistic environments made possible by integrating the AirSim simulator with Google 3D maps models using the Cesium Tiles plugin. Experiments are conducted in this high-realism simulated environment to evaluate the performance of three distinct mapping algorithms: (1) Direct Sparse Odometry (DSO), (2) Stereo DSO (SDSO), and (3) DSO Lite (DSOL). Experimental results evaluate algorithms based on their measured geometric accuracy and computational speed. The results provide valuable insights into the strengths and limitations of each algorithm. Findings quantify compromises in UAV algorithm selection, allowing researchers to find the mapping solution best suited to their application, which often requires a compromise between computational performance and the density and accuracy of geometric map estimates. Results indicate that for UAVs with restrictive computing resources, DSOL is the best option. For systems with payload capacity and modest compute resources, SDSO is the best option. If only one camera is available, DSO is the option to choose for applications that require dense mapping results.
@@ 1 Introduction
UAVs, also known as drones, have transcended conventional applications to become indispensable tools across an array of disciplines, from environmental monitoring and precision agriculture to disaster response and infrastructure inspection. At the heart of their efficacy lies the sophisticated interplay between UAVs and mapping algorithms, which serve as the backbone for converting raw sensor data into coherent, high-fidelity maps. These algorithms play a pivotal role in navigating complex terrains, extracting meaningful information, and ensuring precise localization of the UAV in real time. From traditional photogrammetry to advanced techniques like Simultaneous Localization and Mapping (SLAM), these algorithms continuously evolve to meet the diverse demands of UAV applications ranging from agriculture and forestry to disaster response and urban planning.
@@
Mapping algorithms for UAVs are significantly influenced by flight altitude, dictating the scale of environmental perception and mapping capabilities. For high-altitude flights, the imagery changes between successive frames are slower than for low-altitude flights, which allows more overlap/correspondence between successive frames. However, as altitude increases, challenges such as reduced sensor performance, diminished feature visibility, and heightened geometric distortions emerge. While 3D or 2D laser scanners generate effective terrain models, their weight and sensitivity to ground proximity pose challenges. Compact depth-sensing devices, though commercially available, often fall short in operational range. Camera-based mapping systems, while lightweight and scalable, face accuracy challenges at high altitudes due to reduced texture and discernible features. This limitation hampers feature tracking and matching, impacting overall mapping algorithm performance. High-altitude flights also amplify drift and uncertainty in UAV trajectory estimation, particularly affecting SLAM algorithms relying on sensor fusion. The accumulation of errors over time compromises poses estimations, emphasizing the critical consideration of flight altitude in optimizing mapping algorithm outcomes. This article focuses on multirotor UAVs and analyzes UAV algorithm performance at altitude ranges from 12 m to 20 m from the ground, which is considered to be “low-altitude” in this article. Investigations for this context provide an analysis of key sensor options and their strengths and weaknesses. Specific recommendations are also provided for light-duty UAVs (U.S. military UAS Group 1).
@@
Mapping algorithms tailored for high-speed UAVs address the specific demands of dynamic and rapid flight scenarios. Real-time operation in these contexts is imperative, necessitating synchronization and integration of data from diverse sensors such as LiDAR, cameras, and inertial measurement units (IMUs). Adaptive navigation is equally crucial to accommodate the UAV’s swift maneuvers and maintain mapping precision. Overcoming challenges related to large distances covered between sensor readings during high-speed flights is essential for achieving precise mapping results. Additionally, robustness in the face of environmental variability, including changes in lighting, weather conditions, and terrains, is vital. High-speed UAVs, integral in applications like surveillance and emergency response, benefit from ongoing advancements in mapping algorithms. These improvements enhance effectiveness, allowing UAVs to navigate rapidly changing environments and deliver precise and timely mapping outcomes. This article analyzes multirotor UAV algorithm performance for speed ranges from 15 m/s to 20 m/s which corresponds to the maximum speed for typical commercially available platforms in this Group [kakavitsas2024comparison].
@@
In recent decades, research investigating methods for 3D reconstruction from images has thrived. Examples of approaches include Structure-from-Motion (SfM) algorithms [szeliski2022computer, spetsakis1991multi, triggs2000bundle, lourakis2009sba, crandall2012sfm, meinen2020mapping] and stereo reconstruction (Stereo3D) algorithms [gennery1977stereo, cao2015perception, geiger2011stereoscan, esteban2004silhouette, krutikova2017creation]. These approaches are the algorithms that can be used as components of a SLAM system. SfM and Stereo3D differ in both computation methods and output formats. SfM algorithms analyze a sequence of 2D images from a camera and estimate the relative motion of the camera and the geometric structure of the observed 3D scene. Motion estimates include the camera pose, i.e., position and orientation, at each recorded image and the scene 3D structure observed in each image. Stereo3D estimates 3D scene structure from a pair of 2D images captured simultaneously by two cameras with known relative positions. Depth information is derived from the correspondence of observed scene points between the two images. While both SfM and Stereo3D target 3D scene reconstruction, they excel in different applications and scenarios.
The contributions of this article include:
@@
  • A comprehensive analysis outlining the strengths and limitations of state-of-the-art SfM and stereo reconstruction algorithms;
@@
  • A benchmark of the geometry accuracy and computation speed of various mapping algorithms;
@@
  • A theoretical foundation for sensor selection tailored to low-altitude and high-speed UAV mapping applications;
  • A technical approach for extracting high fidelity geometric models from Cesium Tile data to perform analysis on 3D mapping and odometry algorithms;
@@
  • An innovative approach to simulate realistic flights, utilizing Unreal Engine for high-realism environment synthesis, Cesium plug-in for geographical context, AirSim for vehicle dynamics, and PX4 Autopilot for precise vehicle control.
These contributions provide researchers new insight into how to best adopt mapping technologies for their UAV design in low-altitude and high-speed drone applications.
An initial discussion evaluates the theoretical suitability of a wide variety of sensors for this application and eliminates many sensors from candidacy for various technical reasons. Subsequent evaluation of algorithms is contingent on the proposed selection of best-practice sensors for this context.
Mapping algorithm analysis surveys current state-of-the-art real-time reconstruction algorithms suited to the sensors that were previously identified as appropriate for low-altitude and high-speed multirotor UAV mapping applications. From a wide array of possible algorithms, three were evaluated: (1) Direct Sparse Odometry (DSO) [engel2017direct], (2) Stereo Direct Sparse Odometry (SDSO) [wang2017stereo], and (3) Direct Sparse Odometry Lite (DSOL) [qu2022dsol]. While many algorithms are available in the literature, the selected algorithms provide a representative sampling of reconstruction methods for the recommended camera sensors.
@@ 2 Related Work
This article compares three methods for 3D mapping in terms of their suitability for use on Group 1 UAVs at high-speed, low-altitude flight. Discussion of current sensing options indicates that camera-based methods are well-suited to this application. Experiments use a simulated environment to evaluate leading camera-based methods. For these reasons, a review of the related literature to this article is divided into three parts:
  • A comparison of SfM and stereo3D reconstruction methods including recent leading implementations of these methods;
  • A compact review of three 3D-from-images algorithms;
  • A review of different 3D simulation options for developing and evaluating these mapping algorithms for the context of low-altitude high-speed flight.
A comprehensive literature review motivates the methodology and experimental approach for this article. Specifically, the choice of mapping algorithms analyzed and the simulation environment used was based on a comprehensive review of candidate solutions.
2.1 Structure-from-Motion vs. Stereo Reconstruction
Structure from Motion (SfM) and stereo reconstruction are two leading techniques employed in 3D reconstruction. This subsection describes the principles of both techniques to provide insights into their distinctive attributes and how they relate to high-speed low-altitude mapping applications.
@@ 2.1.1 Structure-from-Motion
SfM [hartley2003multiple] is the process of reconstructing a 3D structure from its projections into a series of images taken from different viewpoints. It leverages the relative movement between a camera and objects in a scene to reconstruct the 3D structure. SfM estimates the camera poses and the spatial arrangement of points in the scene by analyzing the changes in perspective across multiple images. SfM has been extensively studied and applied in diverse fields, including 3D modeling [eltner2020structure, hasheminasab2020gnss], augmented reality [kalacska2021comparing, mooser2009applying], autonomous navigation [mumuni2022bayesian, zhanabatyrova2024structure], and remote sensing [turner2012automated, fujiwara2022comparison]. Researchers have explored various algorithms and optimization methods to enhance the accuracy [szeliski2022computer, caroti2015accuracy, deliry2021accuracy, lindenberger2021pixel] and efficiency [triggs2000bundle, cui2021view, crandall2012sfm, meinen2020mapping] of SfM, making it a robust solution for scenarios where camera poses change dynamically, a common occurrence in high-speed low-altitude flights.
@@ 2.1.2 Stereo Reconstruction
Stereo reconstruction (stereo3D) involves the process of estimating the 3D structure of a scene from a pair of 2D images captured by two cameras with known relative positions. By analyzing the disparities between the two images, stereo reconstruction algorithms can calculate the depth information of the scene points. This depth information allows for the creation of a 3D representation of the scene. Stereo3D is critical to enabling autonomous capabilities in a wide range of fields including robotics [islam2020stereo, pillai2016high], autonomous vehicles [kemsaram2020stereo, cao2015perception], and 3D modeling [liu2022planemvs, shang2023topology, krutikova2017creation].
Figure 1a illustrates the epipolar geometry of two pinhole cameras observing a 3D point $\mathbf{M}$. Stereo reconstruction estimates $\mathbf{M}$’s distance by analyzing its projections $\mathbf{m}$ and $\mathbf{m}^{\prime}$. The baseline $B$ connects camera origins $\mathbf{CL}$ and $\mathbf{CR}$, defining epipolar geometry with epipoles $\mathbf{e}$ and $\mathbf{e}^{\prime}$. The epipolar plane intersects with image planes $\boldsymbol{\pi}$ and $\boldsymbol{\pi}^{\prime}$, forming epipolar lines. According to epipolar geometry, $\mathbf{m}$ in $\boldsymbol{\pi}^{\prime}$ lies on epipolar line $l^{\prime}$. Depth estimation involves finding corresponding points, simplified by image rectification in Figure 1b, ensuring $\mathbf{m}$ and $\mathbf{m}^{\prime}$ align. Depth $d$ is estimated through triangulation represented by Equation (1), considering column differences from $\mathbf{m}$ and $\mathbf{m}^{\prime}$ to the center of the left and right images, baseline $B$, focal length $f$, and pixel width $\delta$ in the rectified image sensor.
$d=\frac{Bf}{\delta\left(x-x^{\prime}\right)}$ (1)
a
b
@@
Figure 2: The dependency between depth estimation accuracy and the baseline of the stereo camera design for a baseline, $B$ of 34 $cm$, based on [irmisch2017camera].
SfM can be computationally intensive and requires feature matching and bundle adjustment for robust results. Stereo3D, with fixed camera positions, is typically less computationally intensive and more straightforward compared to SfM. SfM systems estimate the scene structure to an unknown scale and usually require fusion with other metric data, e.g., from an IMU or a GPS sensor, to make estimated geometric measurements consistent with the real geometric scene structure. Stereo3D directly estimates the scene structure and uses the baseline distance to provide scene scale estimates that are metrically consistent with the 3D scene geometry and do not require sensor fusion to recover the unknown scale.
@@ 2.2 Mapping Algorithms
This article focuses on the following three representative state-of-the-art real-time algorithms to investigate their applications to multirotor UAV-borne mapping:
  • Structure-from-Motion: Direct Sparse Odometry (DSO) [engel2017direct];
  • Stereo Reconstruction: Stereo Direct Sparse Odometry (SDSO) [wang2017stereo] and Direct Sparse Odometry Lite (DSOL) [qu2022dsol].
@@ 2.2.1 DSO: Direct Sparse Odometry
DSO is a visual odometry technique that adapts SfM methods for 3D reconstruction. It directly estimates the camera motion and the sparse 3D structure of the environment from a sequence of 2D images by minimizing photometric errors. DSO differs significantly from traditional techniques by directly optimizing photometric errors in images, without relying on keypoint detectors or geometric priors. For a point, $\mathbf{p}$ in reference frame $I_{i}$, observed as $\mathbf{p^{\prime}}$ in target frame $I_{j}$, the photometric error, given by Equation (2), is formulated as the weighted Sum of Squared Differences (SSD) over a small neighborhood of pixels.
@@
where $\mathcal{N}_{\mathbf{p}}$ is the set of pixels in the SSD; $(t_{i},t_{j})$ the exposure times of the frame $I_{i}$ and $I_{j}$; $(a_{i},b_{i},a_{j},b_{j})$ the brightness transfer variables defined in DSO for frame $I_{i}$ and $I_{j}$, respectively, and $\|\cdot\|_{\gamma}$ is the Huber norm. In addition to using robust Huber penalties, a gradient-dependent weighting $w_{\mathbf{p}}$ is applied. Further, $\mathbf{p}^{\prime}$ stands for the projected point position of $\mathbf{p}$ with inverse depth $d_{\mathbf{p}}$, given by
@@
where $\Pi_{\mathbf{c}}:\mathbb{R}^{3}\rightarrow\Omega$ denotes projection, $\Pi_{c}^{-1}:\Omega\times\mathbb{R}\rightarrow\mathbb{R}^{3}$ denotes back-projection, $\mathbf{c}$ denotes the intrinsic camera parameters, and $\mathbf{T}_{i},\mathbf{T}_{j}\in\textbf{SE}(3)$ are the camera poses represented by transformation matrices for frame $I_{i}$ and $I_{j}$.
To minimize the photometric error between the corresponding points in two frames, DSO incorporates a fully direct probabilistic model that jointly optimizes all model parameters, including camera motion and geometry, represented as inverse depth in a reference frame. The optimization is accomplished using the Gauss-Newton algorithm in a sliding window [leutenegger2015keyframe].
2.2.2 SDSO: Stereo Direct Sparse Odometry
SDSO is a stereo version of DSO. In a monocular mapping system like DSO, to initialize the whole system, i.e., to track the second frame with respect to the initial one using Equation (2), the inverse depth values $d_{\mathbf{p}}$ of the points in the first frame are required. In DSO, the points are initialized to have random depth values ranging from 0 to infinity, corresponding to a large depth variance. Unlike that, SDSO uses stereo matching to estimate a semi-dense depth map for the first frame, which significantly increases the tracking accuracy. The constraints from static stereo introduce scale information into the system. They also provide good geometric priors to temporal multi-view stereo.